OpenStudy (anonymous):

Which expression describes the rule of this sequence? 3, 6, 9, 12, 15, 18

OpenStudy (anonymous):



OpenStudy (anonymous):

or

OpenStudy (anonymous):

3n, 3+3 (3xn) or 3 + n

OpenStudy (angela210793):

as u can see this is an arithmetic progression cause every number is got by adding 3 to the previous number,so d=a2-a1=6-3=3 u can find an(general formula) by using an=a1+d(n-1) a1=3 d=3 substitute and u'll get an=3+(n-1)*3 an=3+3n-3 an=3n